1. Suppose the consumption function is

      C = $309 billion + 0.9Y

If the government increases spending on new goods and services by $50 what is amount of the cumulative shift in AD? Enter your answer in whole numbers. For example, if your answer is $450.25 then enter 450

2. If the AD shortfall is $600 billion and the MPC is 0.8, how large is the desired fiscal stimulus?

If your answer is $50 billion then enter 50
Answer #1

1.Multiplier (k)=1/1-MPC

k=1/1-0.9

k=10

Increase in government stimulus=50

Cumulative change in GDP=k*change in spending

Change in GDP=10*50=500

2.600=k*desired fiscal stimulus

K=1/1-0.8

K=5

Desired fiscal stimulus=600/5=120

Required amount of fiscal stimulus=120